A survey shows that after production 5% of a packet of chips are underweight. If a sample of 20 packets were selected randomly, find the probability that at most 2 packets are underweight.
1
Expert's answer
2020-10-01T06:14:56-0400

solution


From the binomial distribution;


"P(X=r)=\ _nC_r *p^r(1-p)^{n-r}"

Where p is the probability of success, n is the number of trials and r, the number of successes in the trials


Let the Probability of getting underweight chips be be P:


"p =5\%"

The probability that atmost 2 are underweight is the probability of getting 0, 1 or 2 underweight chips.

"p(X \leq 2) = p(X=0)\ + \ p(X=1)\ + \ p(X=2)"


P( X = 0):


"P(X=0)=\ _{20}C_0 *0.05^0(1-0.05)^{20-0}""=0.3585"

P(X = 1):


"P(X=1)=\ _{20}C_1*0.05^1(1-0.05)^{20-1}""=0.3775"

P(X = 2):


"P(X=2)=\ _{20}C_2 *0.05^2(1-0.05)^{20-2}""=0.1887"

Therefore P( X "\leq" 2):


"0.3585+0.3775+0.1887 =0.9245"

answer: the probability of getting at most 2 underweight chips is 0.9245